# Position: Interactive Generative Video as Next-Generation Game Engine

### v1

@@ 4.3 Control
❑ Concept. As shown in Fig. 5, the control module manages user control of the virtual world through two aspects: (1) navigation control enables players to navigate and explore the virtual world through camera and character movement. For example, in racing games, players use arrow keys or “WASD” for acceleration, braking, and steering, while in open-world games, players typically use “WASD” for character movement, mouse for camera rotation, and space bar for jumping or climbing. (2) interaction control allows players to manipulate objects within the virtual environment. For instance, in simulation and management games, players use left mouse clicks to select and place buildings, right clicks to rotate structures, and keyboard shortcuts like ‘E’ to access inventory or ‘Q’ to demolish objects.
@@
• The technical implementation of control mechanisms has been well-studied. Common approaches include: (1) Cross Attention (Yu et al., 2025; Valevski et al., 2024; Feng et al., 2024), where navigation control signals are transformed into conditional features that serve as keys and values, while the video features serve as queries. (2) Another approach uses external Adaptors (Che et al., 2024; Wang et al., 2024b), which directly fuses control features with video features.
• While control is easily mastered in fixed scenes, it should generalize to open-domain scenarios. Some works (Yu et al., 2025; Feng et al., 2024; DeepMind, 2024) have leveraged video generation priors for this purpose, but generalizing complex actions with limited control annotations remains challenging and requires further exploration.
• Learning control, especially for interaction control, goes beyond mechanical execution and requires understanding the underlying rules of how interactions change the environment (as part of physical laws). Following a data-driven approach, future work aims to collect large-scale gaming datasets (Yu et al., 2025; Che et al., 2024) and improve the learning of these interaction rules. Physical laws will be further discussed in the Dynamics module in Sec. 4.5.
@@
Figure 6: The Memory module of IGV consists of two components: static memory and dynamic memory.
@@ 4.4 Memory
Figure 7: The Dynamics module of IGV focuses on two aspects: physical laws and physics tuning.
• Current methods mainly rely on attention-based memory, utilizing attention’s inherent ability to remember historical frames through cross-attention between historical and predicted frames (Valevski et al., 2024; Decart, 2024). However, this approach is unreliable and faces limitations in both precision of memory preservation and limited window size.
• Another promising solution is using dedicated memory structures, which can be implemented either as implicit high-dimensional features (Kim et al., 2020) or explicit 3D representations (Menapace et al., 2022, 2024; Yu et al., 2024a, b). These structures serve as conditional controls for the generation module, ensuring consistent preservation of static elements. While similar approaches have been explored in 3D generation (Yu et al., 2024b; Ma et al., 2024), their adaptation as memory mechanisms for game video generation requires further investigation.
• Dynamic memory, which needs to record animations, actions and trajectories, requires dynamic video datasets, especially those containing significant motions like human movements. Collecting and annotating such high-quality dynamic data remains challenging but crucial for advancing this research direction.
@@ 4.5 Dynamics
Figure 8: The Intelligence module of IGV implements two aspects: causal reasoning and self-evolution.
@@
❑ Technical Approaches and Future Directions.
• A straightforward data-driven approach to implementing physical laws learns probability distributions from large-scale video data to generate physically plausible predictions (OpenAI, 2024; NVIDIA, 2025). However, this requires extensive high-quality videos demonstrating diverse physical phenomena, which remains a significant challenge in current research.
• Alternatively, physics-based memory control offers a more direct approach by simulating physics in the memory space and using it as conditional control for video generation, potentially providing more precise physical behaviors. A specific solution is to use video generation models as renderers on top of physics simulators, which ensures perfect physics compliance. However, this approach is limited to physical phenomena that can be mathematically formulated in physics engines.
• Furthermore, establishing appropriate benchmarks for evaluating the physical accuracy of generated videos represents a crucial research direction (Qin et al., 2024a), as it would help identify limitations and guide improvements in physical dynamics modeling.
• Physics tuning capability, though often overlooked in current research, is crucial for models to truly understand and manipulate physical knowledge. Collecting synthetic video data with annotated physical parameters from simulations could be a viable solution. We raise this challenge to encourage future research efforts in this direction.
4.6 Intelligence
❑ Concept. As Demonstrated in Fig. 8, the Intelligence module implements two key aspects: (1) Reasoning: This capability enables long-term causal inference based on initial conditions, creating deeply immersive virtual worlds. For example, the system can predict how a kingdom’s economy and social structure might evolve over centuries based on its initial resources and policies, or simulate wildlife migration patterns when environmental conditions change, such as animals seeking new water sources after a river dries up. Similar mechanics can be found in grand strategy games like Crusader Kings and ecosystem simulations like Planet Zoo. (2) Self-Evolution: This capability goes beyond merely generating continuous video streams with changing virtual worlds; it enables virtual worlds to continuously develop, evolve, and generate new knowledge, rules, and behaviors through emergent properties. In simulation games, civilizations could naturally emerge and form their own cultures, ecosystems could develop new species, and cities could grow and adapt organically. Such technology could eventually realize a metaverse similar to The Matrix, where countless agents and players live in self-evolving virtual worlds.
❑ Technical Approaches and Future Directions.
• Similar to large language models with their powerful causal reasoning abilities, implementing reasoning capabilities requires video generation models to possess a causal structure. This means new video generation must be conditioned on historical videos through an autoregressive generation approach (as discussed in Generation Module in Sec. 4.2), while also requiring large-scale data training to develop robust causal inference capabilities.
• An alternative approach is to leverage large language models or multimodal large language models for causal reasoning, working in conjunction with video generation models. This strategy appears promising, particularly in developing unified understanding and generation models that can perform both linguistic causal reasoning and visual video generation (Zhou et al., 2024a; Xie et al., 2024). This represents a significant direction for future research.
• Furthermore, if all previously mentioned capabilities including physics understanding, physical simulation, and causal reasoning are successfully implemented and demonstrate powerful performance in the future, we might witness the emergence of remarkable self-evolution capabilities. This convergence of advanced capabilities could potentially lead to truly autonomous virtual worlds, such as metaverses inhabited by countless intelligent agents, or brain-in-a-vat worlds similar to those depicted in The Matrix.
4.7 Gameplay
❑ Concept. The GamePlay Module builds upon Interactive Generated Video (IGV) by implementing external Game Rules, which are designer-imposed rules such as game objectives, rewards, penalties, and constraints that shape the virtual world’s gameplay experience. These include scoring systems in Tetris, health and damage systems in Dark Souls, mission objectives and reward structures in Grand Theft Auto, achievement systems in Minecraft, time limits in Mario, competitive ranking systems in League of Legends, and quest completion rewards in World of Warcraft.
❑ Technical Approaches and Future Directions.
• The implementation of the GamePlay module primarily relies on agent systems empowered by large language models (Achiam et al., 2023; Touvron et al., 2023) or multimodal large models (Liu et al., 2024), leveraging their sophisticated reasoning, planning, and decision-making capabilities. Various gameplay aspects, including level design, difficulty scaling, reward-penalty determination, and NPC character development with dialogue systems, can be implemented through specialized agents.
• Another key research challenge lies in developing a unified multi-agent system framework specifically for game environments. While existing single agents (Li et al., 2023; Wu et al., 2023) show promise, integrating multiple agents to collaboratively control game elements while maintaining coherent gameplay remains an open research problem.
• Another practical research direction is exploring how agents and agent systems can enable dynamic, adaptive game rules, including reward and penalty mechanisms. As players progress through games, their skill levels, capabilities, and experience continuously evolve, making it essential to adaptively adjust difficulty levels and reward-penalty systems accordingly.
5 Levels of Generative Game Engine
We propose a five-level maturity model (L0-L4) to evaluate generative game engines and guide their future development. This framework helps assess current technologies and identify key research directions in generative game engines. Below we detail each level, with an overview of the maturity model presented in Table 1.
Level 0: No AI-Assisted Assets Generation. At this foundational level, game engines rely entirely on manually crafted content without any AI-generated elements. All game assets and rules must be pre-designed by developers during the development phase. Classic examples include Super Mario, where each level layout is carefully hand-crafted, and Tetris, where the game rules and piece designs are fixed. This traditional approach ensures precise control over game content but requires substantial development resources, and players can only experience limited pre-designed content.
Level 1: AI-Assisted Assets Generation. At this level, game development transitions from purely manual processes to AI-assisted creation and integration of game assets and logic, operating both during development and gameplay. During development, AI tools can create diverse game assets to reduce content creation workload. For instance, in Cyberpunk 2077, developers can utilize image generation models like Stable Diffusion (Rombach et al., 2022) to create varied textures for neon billboards, trash piles, and urban details throughout Night City. During gameplay, the game engine can generate short segments in response to player actions, such as unique explosion animations when a player destroys a bridge in an open-world game, or dynamic NPC dialogues in games like AI Dungeon. While this approach accelerates game development and adds variability to common game events, the overall game framework remains pre-designed and still requires significant human intervention and curation.
Level 2: Physics-Compliant Interactive World Generation. This level marks a paradigm shift from traditional manual-centric development to a new paradigm centered on interactive video generation, representing the Next-Gen AI-Driven Generative Game Engines. The game engine continuously generates physics-compliant video content based on player interactions in real-time. For example, when a player sets fire to a wooden bridge, the engine dynamically generates not only the realistic blazing effects but also adapts the game world accordingly, such as rerouting enemy paths around the destroyed structure. While many current research works operate at this level (Valevski et al., 2024; Yu et al., 2025; Decart, 2024), significant improvements are still needed in physics understanding, simulation realism, and the generalization of interaction capabilities.
Level 3: Causal-Reasoning World Simulation. Building upon Level 2’s interactive physics-compliant generation, which focuses primarily on immediate physical responses, this level introduces sophisticated causal reasoning across time in world simulation to address the limitation of short-term effects. The game engine maintains a complex world model that understands both player actions and intrinsic world logic rules, generating video content that reflects long-term cause-and-effect relationships. For example, when a player assassinates a faction leader in Act 1, the engine simulates the resulting political instability, leading to city-wide riots and power struggles that emerge in Act 3. Through this deep understanding of causality, the game creates complex storylines where players’ early choices naturally shape the world’s future development.
Level 4: Self-Evolving World Ecosystem. Building upon the physics-compliant generation of Level 2 and causal reasoning of Level 3, as these capabilities (physics understanding, simulation, and causal reasoning of internal rules) continuously advance, the model emerges with powerful self-evolution abilities. The game world becomes a self-evolving ecosystem where complex systems emerge from initial rules and interactions. For example, as the NPC population grows, they autonomously organize into governance structures and establish trade networks, exhibiting emergent social behaviors beyond their initial programming. At this stage, the generative game engine will be capable of creating virtual worlds similar to those seen in Ready Player One or The Matrix, where future players can not only play but potentially live within these worlds. This advancement will not only revolutionize the gaming industry but also profoundly impact various aspects of human society.
6 Alternative Views
Alternative Views #1: Current video generation models lack sufficient physical understanding, making it challenging to implement GGE. Recent benchmarks (Qin et al., 2024a; Kang et al., 2024; Motamed et al., 2025) evaluating video models’ grasp of physical laws have consistently revealed major limitations. These studies, despite using different evaluation methods, show that current models struggle to accurately predict object trajectories, velocities, and interactions. This gap between visual realism and physical understanding remains a significant hurdle for game development.
Potential Solution #1: WorldSimBench (Qin et al., 2024a) improves video models’ physical understanding through the HF-Embodied dataset, while Physics-IQ-benchmark (Motamed et al., 2025) demonstrates that training on large-scale datasets with rich physical interactions can help models learn basic physical principles through observation.
Alternative Views #2: IGV’s current limitations in scene consistency prevent it from serving as GGE’s core. Some approaches (Decart, 2024) show that simple camera movements, like looking up and down, can cause dramatic scene changes (e.g., plains suddenly becoming desert), breaking spatial continuity and gameplay immersion.
Potential Solution #2: Scene consistency can be improved through 3D control mechanisms. Recent works (Ma et al., 2024; Yu et al., 2024b; Xie et al., 2025) demonstrate promising results by reconstructing scene point clouds from generated videos and using them to guide subsequent video generation, enhancing spatial coherence.
Alternative Views #3: IGV’s limited logical reasoning capabilities make it inadequate as GGE’s core. Games require complex logical structures, particularly in narrative progression, which current video generation models cannot effectively handle. Existing IGV approaches (Decart, 2024; Yu et al., 2025; Valevski et al., 2024) only work in simplified settings, acting more as rendering engines rather than addressing the deeper logical demands of game environments.
Potential Solution #3: An approach is to combine video models with large language models (LLMs). While video models alone struggle with logical reasoning, integrating them with LLMs offers a viable solution. Recent unified models (Zhou et al., 2024a) that combine diffusion-based video generation with language understanding show encouraging results in handling both game logic and video content. As this technology matures, IGV could develop the necessary capabilities to function effectively as GGE’s core.
While achieving GGE through IGV faces challenges, targeted model designs and high-quality data collection can address these issues. We believe that with continued advancements in video generation models, IGV can ultimately serve as the core technology for GGE.
7 Conclusion
Impact Statement
This paper proposes Interactive Generative Videos (IGV) as a core technology for the next-generation game engine, Generative Game Engine (GGE). IGV will transform the way generative AI is utilized, further advancing the development of interactive applications. GGE will revolutionize the gaming industry by enabling rapid production of high-quality, unlimited, and adaptive game content.

### v2

@@ 4.2 Control
❑ Concept. As shown in Fig. 5, the control module manages user control of the virtual world through two aspects: (1) navigation control enables players to navigate and explore the virtual world through camera and character movement. For example, in racing games, players use arrow keys or “WASD” for acceleration, braking, and steering, while in open-world games, players typically use “WASD” for character movement, mouse for camera rotation, and space bar for jumping or climbing. (2) interaction control allows players to manipulate objects within the virtual environment. For instance, in construction games, players use left mouse clicks to select and place buildings, right clicks to rotate structures, and keyboard shortcuts like ‘E’ to access inventory or ‘Q’ to demolish objects.
@@
The technical implementation of control mechanisms has been well-studied. Common approaches include: (1) Cross Attention gamefactory ; gamengine ; matrix , where control signals are transformed into conditional features that serve as keys and values, while the video features serve as queries. (2) Another approach uses external Adaptors gamegenx ; motionctrl , which directly fuses control features with video features.
While control is easily mastered in fixed scenes, it should generalize to open-domain scenarios. Some works gamefactory ; matrix ; genie2 have leveraged video generation priors for this purpose, but generalizing complex actions with limited control annotations remains challenging and requires further exploration.
Learning control, especially for interaction control, goes beyond mechanical execution and requires understanding the underlying rules of how interactions change the environment (as part of physical laws). Following a data-driven approach, future work aims to collect large-scale datasets gamefactory ; gamegenx and improve the learning of these interaction rules. Physical laws will be further discussed in Sec. 4.4.
@@ 4.3 Memory
Figure 6: The Memory module consists of static and dynamic memory.
@@
Current methods mainly rely on attention-based memory, utilizing attention’s inherent ability to remember historical frames through cross-attention between historical and predicted frames gamengine ; oasis ; worldmem . However, this approach is unreliable and faces limitations in both precision of memory preservation and limited window size. Another promising solution is using dedicated memory structures, which can be implemented either as implicit high-dimensional features gamegan or explicit 3D representations pe ; pgm ; wonderjourney ; viewcrafter ; see3d ; gen3c . These structures serve as conditional controls for the generation module, ensuring consistent preservation of static elements. The adaptation of these methods as memory mechanisms for game video generation requires further investigation.
@@ 4.4 Dynamics
Figure 7: The Dynamics module focuses on physical laws and physics tuning.
❑ Technical Approaches and Future Directions.
A data-driven approach learns physical laws from large-scale video data sora ; cosmos , though this requires extensive high-quality videos demonstrating diverse physical phenomena zhao2025synthetic . Physics-based memory control offers an alternative by using video generation models as renderers on top of physics simulators physgen ; diffusion-as-shader , ensuring perfect physics compliance but limited to mathematically formulated phenomena. Establishing appropriate benchmarks for evaluating physical accuracy remains crucial worldsimbench ; worldmodelbench to identify limitations and guide improvements. Physics tuning capability, often overlooked in current research, is crucial for models to truly understand and manipulate physical knowledge, and we encourage future research to explore synthetic data with annotated physical parameters as a potential solution.
4.5 Intelligence
Figure 8: The Intelligence module implements reasoning and self-evolution.
❑ Concept. As Demonstrated in Fig. 8, the Intelligence module implements two key aspects: (1) Reasoning: This capability enables long-term causal inference based on initial conditions, creating immersive virtual worlds. For example, the system can predict how a kingdom’s economy and social structure might evolve over centuries based on its initial resources and policies, or simulate wildlife migration patterns when environmental conditions change, such as animals seeking new water sources after a river dries up. Similar mechanics can be found in strategy games like Crusader Kings and ecosystem simulations like Planet Zoo. (2) Self-Evolution: This capability goes beyond generating continuous video streams with changing virtual worlds; it enables virtual worlds to continuously develop, evolve, and generate new knowledge, rules, and behaviors through emergent properties. In simulation games, civilizations could naturally emerge and form their own cultures, ecosystems could develop new species, and cities could grow and adapt organically. Such technology could eventually realize a metaverse similar to The Matrix, where countless agents and players live in self-evolving virtual worlds.
❑ Technical Approaches and Future Directions.
Implementing reasoning capabilities requires video generation models to have a causal structure through autoregressive generation (as discussed in Generation Module in Sec.4.1) and large-scale long-context training far , similar to large language models. Alternatively, leveraging (multimodal) large language models for causal reasoning alongside video generation models shows promise for unified understanding and generationzhou2024transfusion ; xie2024show . Furthermore, if all previously mentioned capabilities including physics understanding, physical simulation, and causal reasoning are successfully implemented and demonstrate powerful performance in the future, we might witness the emergence of remarkable self-evolution capabilities. This convergence of advanced capabilities could potentially lead to truly autonomous virtual worlds, such as metaverses inhabited by countless intelligent agents, or brain-in-a-vat worlds similar to those depicted in The Matrix.
4.6 Gameplay
❑ Concept. The GamePlay Module builds upon IGV by implementing external Game Rules, which are designer-imposed rules such as game objectives, rewards, penalties, and constraints that shape the virtual world’s gameplay experience. These include scoring systems in Tetris, health and damage systems in Dark Souls, mission objectives and reward structures in Grand Theft Auto, achievement systems in Minecraft, time limits in Mario, competitive ranking systems in League of Legends, and quest completion rewards in World of Warcraft.
❑ Technical Approaches and Future Directions.
The implementation of the GamePlay module primarily relies on agent systems empowered by large language models gpt4 ; llama or multimodal large models llava , enabling various gameplay aspects including level design, difficulty scaling, and NPC development. While existing single agents camel ; autogen show promise, key research challenges remain in developing unified multi-agent frameworks for game environments. Another practical research direction is exploring how agents and agent systems can enable dynamic, adaptive game rules, including reward and penalty mechanisms. As players progress through games, their skill levels, capabilities, and experience continuously evolve, making it essential to adaptively adjust difficulty levels and reward-penalty systems accordingly.
5 Levels of Generative Game Engine
We propose a five-level maturity model (L0-L4) to evaluate GGE and guide their future development. This framework helps assess current technologies and identify key research directions in GGE. Below we detail each level, with the overview table presented in Appendix B.
Level 0: No AI-Assisted Assets Generation. At this foundational level, game engines rely entirely on manually crafted content without any AI-generated elements. All game assets and rules must be pre-designed during the development phase. Classic examples include Super Mario, where each level layout is carefully hand-crafted, and Tetris, where the game rules and piece designs are fixed. This approach enables precise control but requires heavy resources and restricts players to fixed content.
Level 1: AI-Assisted Assets Generation. Game development combines manual processes with AI-assisted creation of assets and logic during development and gameplay. AI tools generate diverse assets to reduce content creation workload. For instance, in Cyberpunk 2077, developers can utilize image generation models like Stable Diffusion sd to create varied textures for neon billboards, trash piles, and urban details throughout Night City. During gameplay, the engine generates segments, such as unique explosion animations when a player destroys a bridge in an open-world game, or dynamic NPC dialogues in games like AI Dungeon. While this approach speeds up development and adds variety, the framework remains pre-designed and needs significant human intervention and curation.
Level 2: Physics-Compliant Interactive World Generation. This level shifts from manual-centric development to interactive video generation, representing AI-Driven Generative Game Engines. The engine continuously generates physics-compliant content based on player interactions in real-time. For example, when a player sets fire to a wooden bridge, the engine dynamically generates not only the realistic blazing effects but also adapts the game world accordingly, such as rerouting enemy paths around the destroyed structure. While many works operate at this level gamengine ; gamefactory ; oasis , significant improvements are needed in physics understanding, simulation realism, and generalizable interaction.
Level 3: Causal-Reasoning World Simulation. Building on Level 2’s physics-compliant generation, which focuses on immediate responses, this level adds causal reasoning across time to address short-term limitations. The engine maintains a world model that understands player actions and logic rules, generating content that reflects long-term cause-and-effect relationships. For example, when a player assassinates a faction leader in Act 1, the engine simulates the resulting political instability, leading to city-wide riots and power struggles that emerge in Act 3. Through this understanding, the game creates storylines where players’ early choices shape the world’s future development.
Level 4: Self-Evolving World Ecosystem. Building on Level 2’s physics generation and Level 3’s causal reasoning, as these capabilities continuously advance, the model emerges with self-evolution abilities. The game world becomes a self-evolving ecosystem where complex systems emerge from initial rules and interactions. For example, as the NPC population grows, they autonomously organize into governance structures and establish trade networks, exhibiting emergent social behaviors beyond their initial programming. At this stage, the engine will create virtual worlds similar to those in Ready Player One or The Matrix, where players can not only play but potentially live within these worlds. This advancement will revolutionize gaming and profoundly impact human society.
6 Alternative Views
Alternative View #1: While GGE represents an automated approach to game content generation, it is worth examining whether it shares the same limitations as Procedural Content Generation (PCG) methods, specifically the tendency to produce repetitive content and the presence of difficult-to-fix bugs that potentially limit their practical applications.
Potential Solution #1: GGE differs fundamentally from Procedural Content Generation (PCG). PCG creates infinite content by randomly combining limited assets and predefined logic rules. In contrast, GGEs learn from massive datasets, acquiring knowledge of unlimited assets and world logic rules. Unlike PCG’s meaninglessly repetitive content generation, GGE can create truly diverse content, similar to how AI image generation has enabled diverse, high-quality artworks on Civitai civitai . Additionally, GGEs implicitly model logic rules and leverage control modules for precise control, avoiding PCG limitations such as difficult control, procedural bugs and debugging needs.
Alternative View #2: Given that traditional rendering pipelines in standard game engines offer efficient asset rendering and allow more resources to be allocated to gameplay enhancement, why should we adopt IGV instead of maintaining the traditional approach that prioritizes gameplay dynamics over graphical realism?
Potential Solution #2: Traditional rendering pipelines efficiently handle graphics, allowing developers to focus more resources on gameplay rather than graphics. This raises concerns that IGV might shift too many resources toward visual realism at the expense of gameplay quality, a trade-off that many developers would find unreasonable. However, IGV represents a paradigm shift that enhances both graphical realism and gameplay quality together, rather than trading off one for the other. We analyze this from four aspects: (1) Complete system: IGV is not just a generation model, but a system integrating Control, Memory, Dynamics and Intelligence. Beyond improving graphics, it enhances gameplay through, for example, personalized game content, infinite explorable experiences, and intelligent NPC behaviors. (2) Enhanced gaming experience: IGV enables dynamic, customized, and infinitely explorable experiences, which traditional game development with triangle rendering and added gameplay logic cannot easily achieve. (3) Enhanced creative freedom: IGV’s vitual world generation capabilities free developers from focusing on graphics, allowing them to concentrate on gameplay design. Its controllability enables developers to freely exercise creativity in designing more innovative gameplay experiences. (4) Positive industry impact: IGV’s efficiency and capabilities accelerate game development and lower entry barriers. This attracts more developers, resulting in more creative games and enriching the gaming experience industry-wide.
We also address an additional alternative view regarding the concerns about GGE costs in Appendix C.
7 Ethical Issues
Several key ethical issues need to be carefully considered in the development and application of GGE: copyright concerns (determining ownership and protection of AI-generated content), security issues (preventing the generation of harmful content), creativity concerns (whether AI enhances or limits human creative expression), democratization implications (the impact of lowering barriers to game creation), and labor concerns (potential effects on gaming industry workers). These critical issues require thorough discussion and resolution, which we address in detail in the Appendix D.
8 Conclusion
Appendix A Preliminaries
A.1 Video Generation Models
Video Generation Models aim to synthesize temporally consistent video sequences $\textbf{x}=\{x_{0},x_{1},\dots,x_{t}\}$, where $x_{i}$ indicates the $i$-th frame. The video is generated by an autoregressive model [81, 110, 53, 102], a diffusion model [40, 77], or a masked transformer model [11, 122]. With the rise of diffusion models [41, 91, 90, 61, 64], which have now become the mainstream approach in video generation due to its high-quality generation performance, significant progress has been achieved in this field [77, 117, 52, 22, 87, 94, 85, 2, 8, 99, 54].
Conditional Video Generation is formulated as $p(\textbf{x}|c)$ and it varies depending on the type of control signal $c$ which denotes conditions such as text prompts or other control signal. Approaches such as [36, 72, 109] incorporate images as control signals for the video generator, improving both video quality and temporal relationship modeling. Methods such as Direct-a-Video [116], MotionCtrl [104], and CameraCtrl [38] use camera embedders to adjust camera poses, enabling control over camera movements in generated videos. 3DTrajMaster [28] extends this capability by transforming 2D camera signals into 3D for more advanced control. ReCamMaster [5] re-shoots the source video with novel camera trajectories.
Autoregressive Video Generation Models. Since game videos require variable-length or even infinite-length generation to enable interactive game experience, autoregressive mechanism is necessary. Autoregressive video generation refers to a process where new frames are generated based on previously generated frames, which can be expressed as $p(x_{1},x_{2},...,x_{n})=\prod_{i-1}^{n}p(x_{i}|x_{1},x_{2},...,x_{i-1})$, where $x_{i}$ indicates the $i$-th frame. An intuitive approach is to adopt GPT-like next-token prediction methods [53, 24, 102]; however, this approach often falls short in terms of generation quality. Relying on the exceptional performance of diffusion models, Diffusion Forcing [14, 89] implements autoregression by applying different levels of noise to different frames, allowing the denoising of new frames with higher noise levels while conditioning on previous frames with lower noise levels. Methods [121, 27, 98, 20] leveraging Diffusion Forcing has achieved remarkable results.
A.2 AI-driven Game Applications
Game Video Generation. Previous works have utilized GANs [51, 50, 67] to generate game videos or used NeRF to reconstruct 3D scenes to simulate the game process [66, 68], but often fall short in terms of generation quality. Using the powerful generative capabilities of diffusion models, some works [3, 98, 114, 20, 27] have produced high-quality game videos. However, the content is typically confined to specific preexisting games. Open-domain Methods [21, 27, 121], by utilizing multi-stage training or large-scale training datasets, create new content for video games.
Game Design Assistant. AI-powered design assistants offer numerous advantages throughout the creative process, depending on the tool, the type of AI and the creative workflow. These systems can streamline development, reduce costs, reduce manual effort, improve team collaboration, and even inspire creativity [58]. In the gaming domain, most existing AI-driven design tools primarily assist by auto-completing an ongoing design [88] or generating multiple design suggestions for creators to evaluate [59, 70, 12, 96].
Intelligent Game Agent. Reinforcement learning has long been the predominant approach in this domain. Early efforts explored the use of hierarchical RL [10, 46, 26, 60, 74, 127, 4] in the context of MineRL competitions [37]. However, due to the absence of guidance from prior knowledge, such approaches often struggle to perform effectively on long-horizon tasks. With the advancement of LLM [1, 97], leveraging their prior knowledge to plan long-horizon tasks has shown promising results. Recent advancements in LLM-related research [100, 105, 128, 80, 111, 44] have significantly propelled the progress of agents in long-horizon tasks.
Appendix B Overview Table for Levels of GGE
In Table 1, we demonstrate the overview of different maturity levels for GGE.
@@ Appendix B Overview Table for Levels of GGE
Appendix C Additional Alternative Views
Alternative View #3: The economic costs of GGE appear to be significant. For instance, the computational overhead is substantial since GGE relies on IGV and LLMs, which are computationally intensive large models. Do these costs prevent IGV-centered GGE from becoming the next generation of game engines? Are these costs we incur for implementing GGE justified by the benefits it brings?
Potential Solution #3: Regarding the computational costs, we believe these can be effectively reduced through technological advancements. Recent works have demonstrated promising advances in efficient autoregressive video generation. On the algorithmic front, CausVid [119] achieves real-time frame generation through distribution matching distillation (DMD) [118], while Cosmos [73] enables real-time generation by combining Medusa speculative decoding, key-value caching, tensor parallelism, and low-resolution adaptation. Additionally, hardware optimizations like GPU parallelization, quantization, and knowledge distillation have significantly accelerated inference speeds for autoregressive models. With ongoing research in efficient models, we believe autoregressive video generation will eventually achieve real-time performance on commonly available hardware accessible to game developers.
Beyond computational costs, other economic considerations include:
  • Data Collection Costs: These will be mitigated as more open-source datasets like GameGen-X [13] become available. While initial training incurs costs, trained models reduce future asset production costs, leading to overall savings.
  • Licensing Costs: Generative models will lower the barrier for developers to create their own new IPs. Building a mutually beneficial ecosystem between developers and gaming companies is also advantageous.
  • Safety Control Costs: While this affects all generative AI, not just IGV, the benefits of incorporating generative AI outweigh these costs, as demonstrated by successful products like Runway [85], Midjourney [69], and ChatGPT [76].
We believe that these costs will not impede GGE’s development or future potential. The technology is continuously evolving, with costs decreasing while model capabilities become increasingly powerful, making the benefits more and more significant. This mirrors the trajectory of large language models. Compared to ChatGPT [76] released in 2022, today’s LLMs demonstrate stronger performance (like DeepSeek-R1 [32]’s reasoning capabilities and GPT-4o’s multimodal generation abilities [75]) while becoming cheaper and more accessible (open-source models like DeepSeek [32] and Qwen [6] now offer performance comparable to commercial models).
While GGE currently faces some cost-related concerns in the short term, these challenges are outweighed by its transformative value. As discussed in Alternative View #2 in Sec. 6, GGE offers significant advantages over traditional game engines, such as personalized gaming experiences, infinitely generated game content, and lowering the barrier to game development so that everyone can become a game designer. These compelling benefits, which are unattainable with traditional game engines, make GGE’s economic costs worthwhile to address and overcome.
Appendix D Ethical Issues
Copyright Issues: How should we determine copyright ownership and protect legitimate copyright interests of all parties involved in GGE-generated games?
Copyright protection presents a new challenge in generative AI development, including IGV, which requires significant attention from both technical and legal perspectives. While this is a complex issue, the industry is actively working towards solutions that can foster the mutual development of AI technology and copyright protection. To address these challenges, we propose several approaches:
Training data for IGV should prioritize the use of copyright-free or properly licensed data sources to minimize legal risks. Game developers can build mutually beneficial partnerships with copyright holders to legally obtain data and share the copyright of the created content. For instance, while noting the Studio Ghibli’s recent copyright concerns with OpenAI, we observe that Ghibli has successful experiences in collaborating with game companies (e.g., development of game "Ni no Kuni"). Such examples demonstrate the feasibility and value of proper copyright collaboration.
Technically, research works [34, 33] on dataset copyright protection and detection of unauthorized training data usage are progressing, which has positive implications for IGV in game development.
Security Issues: What measures can be implemented to prevent the generation of harmful content by generative models such as IGV?
IGV systems are built upon existing video generation models, thus inheriting their established safety measures. Current commercial video generation services like Runway and Sora have implemented comprehensive safety systems that filter out inappropriate content including violence, pornography, hate speech, and other harmful materials. From a technical perspective, safety measures can be implemented through various approaches: (1) Value alignment [19, 86] through techniques like RLHF during the model post-training phase would establish fundamental safety boundaries. This alignment with human preferences and values can effectively constrain the model’s output content; (2) Real-time harmful content detection [103, 39] using VLMs can quickly analyze generated content, identify potential harmful elements, and block inappropriate content in real-time, which is particularly crucial in interactive gaming environments.
Creativity Issues: Can IGV serve as a creative tool, allowing for deep human creativity?
[…]
Democratization Issues: Does democratizing game creation diminish its value?
We believe that democratizing game creation will not diminish its value, but rather enhance the overall value and creativity of the entire field. Here is our analysis and examples:
The democratization of gaming won’t diminish creative value. The widespread availability of technology enables more people to enter this field, generating more diverse creative thinking and innovative designs. Creation value lies in innovation and personalization, not just technical difficulty. Through this technology, even ordinary users can create games with unique characteristics and personal style, which holds its own distinctive value. A good example is the opening up of image generation models, which hasn’t diminished the value of artistic creation. People with varying levels of professional expertise have shared numerous new artistic works on Civitai [18], which has actually enhanced the creativity in this field and the value of its works.
Labor Issues: How should we view the potential negative impact of highly automated productivity tools like GGE on labor in the gaming industry?
We acknowledge the labor impact concern with generative AI. IGV aims to enhance productivity and creativity rather than replace human workers. We advocate for measures like education and support programs to help industry professionals leverage AI tools, ensuring positive industry transformation.
Appendix E Workflow Integration with GGE
It is important to emphasize that the introduction of generative game engines (GGE) will not lead to a single, rigid game development workflow. We provide below a framework example of how GGE can be incorporated into game development workflows
Phase #1: Pre-production Phase • Concept Design: Define core game elements (gameplay mechanics, story, target audience, art style) through LLM consultation and convert to IGV condition prompts. • Prototype Development: Select suitable base models based on computing power and performance requirements, and develop prototypes using initial prompts for feasibility testing.
Phase #2: Production Phase • Asset and Logic Requirements: Create detailed prompts for specific assets and logic requirements (e.g., character model descriptions, area map sketches, level-up rule systems). • Training Data Collection and Model Fine-tuning: Fine-tune models with targeted game data (e.g., collecting copyright-free space movie/game assets for a space exploration game).
Phase #3: Testing Phase • Functionality Testing: Test prompt-based content generation and screen for harmful content. • Compatibility and Performance Testing: Optimize performance across different devices with necessary algorithm/hardware acceleration.
Phase #4: Post-launch Maintenance • Content Updates: Update model parameters and prompts for new content (DLCs, characters, events). • Data Analysis and Optimization: Use player behavior data (with consent) for model fine-tuning and reinforcement learning.
Feasibility Requirements The successful implementation of this workflow relies on these key factors: • (1) Model Capability: Robust base IGV models that support efficient control, fine-tuning, and fast inference. • (2) Data Accessibility: Well-established data sharing and copyright mechanisms that enable legal and cost-effective access to high-quality training data. • (3) Computing Resources: Accessible AI computing infrastructure, either through local hardware resources or affordable cloud computing services.